# Temperature dependence of coercivity for isolated Ni nanowires unraveled by high-sensitivity micromagnetometry

### v1

@@ Abstract
Magnetic nanowires are critical components in fields such as data storage and spintronics, where precise control of their magnetic properties is essential for device optimization. In particular, the behavior of isolated nanowires is often different from that of an ensemble, offering an opportunity to explore the role that dipolar and magnetoelastic interactions play in the latter system. Unfortunately, the comparison between a collection of nanowires and single ones is often poorly characterized, as measuring individual nanowires with weak magnetic signals is a challenging task. In this work, we employ a highly-sensitive micromechanical torsional oscillator to measure the magnetic response of few individual Ni nanowires with $(72\pm 5)$ nm average diameter, fabricated by electrodeposition in anodic aluminum oxide templates as an array and subsequently released from this membrane. When comparing the magnetic properties as a function of temperature between single nanowires and the array, we show that coercivity values of individual nanowires are at least twice as large as for the array in the range $5$ K - $200$ K. Also, we characterize the differences in the hysteresis loops, which are more squared for isolated nanowires, with a high magnetic remanence close to $80\%$ of the saturation value. Our results highlight the crucial role of dipolar and mechanical interactions in modifying the magnetic behavior of nanowires arrays, providing valuable insights for the design and application of nanowires-based magnetic devices.
@@ I Introduction
Using the electrodeposition technique, ferromagnetic Ni NWs with a variety of sizes and morphologies can be synthesized inside the pores of an anodic aluminum oxide (AAO) template Razeeb2009 ; Proenca2012 . The magnetic hysteresis properties of these structures have been extensively studied Whitney1993 ; Sellmyer2001 ; Kato2004 ; Kumar2006 ; Meneses2015 ; Meneses2018 ; Campos2020 . In general, within a wide range of temperatures $T$ between $5$ K and $300$ K, a monotonic reduction in coercivity $H_{c}$ with increasing $T$ is observed in arrays of Ni NWs with diameters larger than $d\approx 90$ nm which are still embedded in the AAO template Adeela2015 . This is a typical behavior also found in bulk multidomain Ni ferromagnets Hayashi1964 , single-domain Ni nanoparticles Fonseca2003 , and Ni films Hanson1997 . Interestingly, the coercivity of thinner NWs still embedded in the template shows an infrequent behavior: $H_{c}$ increases with increasing $T$ and in some cases it exhibits a maximum around room temperature Kumar2006 ; Meneses2018 ; Campos2020 ; Zeng2002 ; Navas2008 ; Michel2017 .
Previous studies suggest that the internal microstructure of each wire (i. e. the mean grain size and the magnetocrystalline anisotropy) and the dipolar interactions between the NWs, are not the origin of this unexpected behavior in $H_{c}$. Instead, this phenomenon could simply be due to the competition between the shape anisotropy of individual NWs; the magnetoelastic anisotropy arising from the mechanical interaction between the NWs, the AAO template, and the Al substrate; and the dipolar interactions between NWs Kumar2006 ; Meneses2018 ; Zeng2002 . Experimental data seem to confirm these conjectures.
In this work, we use a silicon micromechanical torsional oscillator to study the hysteresis behavior of isolated Ni NWs in a temperature range of $5$ K to $200$ K. In absence of the template, substrate and dipolar interactions, we measure a coercivity which depends only on the intrinsic characteristics of the NWs.
The outline of the paper is as follows. In Sec. II, we briefly describe how the Ni NWs were synthesized and released from the template, focusing on the most relevant characteristics of the magnetic sample we have studied. We also describe the basic operation of the micromechanical torsional oscillator. In Sec. III we present the results and discuss their implications. Finally, Sec.IV is devoted to the summary and conclusions.
@@ II Experiments
II.1 Ni nanowires
We study Ni NWs that were synthesized in a previous work Meneses2018 . In particular, we focus on sample DC65 where the NWs were grown by direct current (DC) electrodeposition inside the pores of an AAO template. In Ref. Meneses2018 , the magnetic hysteresis loops of this sample were measured in a Super Conducting Quantum Interference Device (SQUID), in a temperature range of $5$ K to $300$ K. The coercivity of sample DC65 increases with increasing $T$ and exhibits a maximum around room temperature. A priori, it is not clear whether this phenomenon reflects the magnetic properties of a single NW, or rather the collective behavior of an array of NWs still embedded in the AAO template.
Figure 1: (a) SEM image of the silicon micromechanical oscillator showing the NWs stuck to its plate. This structure is attached to two springs which are in turn anchored to the substrate, allowing torsional oscillations forced by an alternating voltage signal connected to the electrodes. When a magnetic field $\mathbf{H}$ is applied, the resonance frequency of the device changes allowing to infer the magnetic moment of the sample. (b) SEM image of a typical Ni NW.
@@ II.1 Ni nanowires
Figure 2: SEM image showing a bundle made up of NWs of approximately $l=9$ $\mu$m in length.
More detailed SEM images show that some of the fragments (supposedly individual NWs) that were glued on the microsensor, are actually bundles formed from just a few NWs, as displayed in Fig. 2. Since these clusters are made up of few NWs, the dipolar interactions between them are extremely low. To estimate the order of magnitude of this unwanted contribution, we roughly calculate the longitudinal component of the magnetic field $H_{z}$ produced by a NW at a transverse distance $r$ from its center of mass, describing this object by two magnetic charges separated by a distance $l$ (the length of the NW) Kumar2006 ; Bertotti1998 :
@@
Here, $m=VM_{b}$ is the magnetic moment of a NW where $V$ and $M_{b}$ are, respectively, its volume and saturation magnetization. In this way, a single Ni NW with $d=72$ nm, $l=9$ $\mu$m, and $\mu_{0}M_{b}=0.61$ T Bertotti1998 , where $\mu_{0}$ is the vacuum permeability constant, produces a field of approximately $\mu_{0}H_{z}\approx 20$ $\mu$T at a distance $r=d$, which is smaller than the Earth’s magnetic field Cullity . This shows that the magnetic behavior of one NW belonging to a bundle as shown in Fig. 2, will not be affected by the presence of its neighbors. This is true only if the number of nanowires involved is very small. In contrast, in a dense array of NWs, the long-range dipolar interaction is significant and therefore cannot be ignored.
In our setup, we glued 11 NWs onto the plate of the micromechanical oscillator, and used SEM images to characterize their geometry. Since the NWs are branched at many points, locally enlarging their cross-sections, we measured an effective diameter of $(d=72\pm 5)$ nm, which is slightly larger than that of the idealized cylindrical nanowire. The collection of NWs features lengths ranging from $5$ to $9$ $\mu$m, resulting in a total sample volume of $V_{n}=(3.2\pm 0.4)\times 10^{-19}$ m3.
@@ II.2 Micromagnetometer
The silicon micromechanical torsional oscillator was manufactured in the MEMSCAP Inc. Foundry MEMSCAP . Similar microdevices have been previously used as micromagnetometers of high sensitivity to measure the hysteresis behavior of manganite nanotubes Dolz2008 ; Antonio2010 ; Dolz2020 as well as mesoscopic samples of a high-$T_{c}$ superconductor Dolz2007 ; Dolz2010 . In simple terms, its operation can be understood as follows. The whole system (the micromechanical oscillator with the NWs stuck on its plate) is cooled under vacuum inside a helium closed-cycle cryogenerator until reaching a given temperature. Then, the microdevice is actuated electrostatically by means of a function generator and its movement is sensed capacitively using a lock-in amplifier. In this way, the resonance frequency of the system is accurately measured which depends, among other things, on the elastic constant of its serpentine springs. When a uniform magnetic field (provided by an electromagnet) is applied along the easy axis of the NWs [see Fig. 1 (a)], the plate of the oscillator experiences an additional torque and therefore the resonance frequency changes. Although extremely small, these shifts in resonance frequency (less than 1 Hz) can be measured and used to infer the magnetization state of the sample (see below). Additional details of the experimental setup can be found in Refs. Dolz2008 ; Antonio2010 .
@@
Figure 3: Resonance curves obtained when $\mathbf{H}$ or $\mathbf{M}$ (or both) are zero and therefore there is no additional magnetic torque (cases I and II), and when these fields are parallel (case III) or antiparallel (case IV) which results in the appearance of a restoring or anti-restoring magnetic torque, respectively.
@@
Typical resonance curves are plotted in Fig. 3. With red squares, we show the amplitude of oscillation as a function of frequency when there is no additional magnetic torque on the system. This happens when either no external field $\mathbf{H}$ is applied, or when the magnetization $\mathbf{M}$ of the sample is zero, which correspond, respectively, to cases I and II drawn in Fig. 4 (furthermore, there is no magnetic torque in the trivial case where both vectors are zero). In particular, when performing a hysteresis experiment, case II arises at the coercive field, $\mathbf{H}_{c}$, since in this instance $\mathbf{M}=0$.
@@
Figure 4: Sketches corresponding to cases I, II, III, and IV. No additional magnetic torque arises when (a) $\mathbf{H}=0$ or (b) $\mathbf{M}=0$ However, when both fields are non-zero, (c) a restoring torque appears if $\mathbf{H}$ and $\mathbf{M}$ are parallel, while (d) this torque is anti-restoring if $\mathbf{H}$ and $\mathbf{M}$ are antiparallel.
@@
A different situation appears when both $\mathbf{H}$ and $\mathbf{M}$ are non-zero. As shown in case III in Fig. 4, if these vectors are parallel any slight twist with respect to the equilibrium position results in the appearance of a restoring magnetic torque, which adds to the torque produced by the serpentine springs. As a consequence, the resonance frequency must increase (note that in any mechanical oscillator, as the restoring force or torque increases, the resonant frequency also increases). The corresponding curve in Fig. 3 (blue circles) shows this positive shift in frequency. On the contrary, if $\mathbf{H}$ and $\mathbf{M}$ are antiparallel as in case IV, the sample experiences an anti-restoring magnetic torque, which reduces the total torque applied to the oscillator’s plate and then the resonance frequency decreases, see the corresponding resonant curve in Fig. 3 (magenta triangles).
@@
The above qualitative description can be transformed into a quantitative formalism by making a series of simple analytical calculations and well-founded approximations. Technical details are given in the Appendix. Such calculations lead to a simple expression that allows the magnetization of the sample to be estimated from the changes in the resonance frequency, $\Delta\nu$, measured in our experiments:
@@
where $\nu_{0}$ is the natural resonance frequency of the micromechanical torsional oscillator (when the magnetic torque is zero), $I$ is the moment of inertia of the whole system along its center rotational axis and, as we defined above, $V_{n}$ is the volume of the magnetic sample.
@@ III Results and discussion
Figure 5 (a) shows the hysteresis loops of $\Delta\nu$ vs $H$ measured at temperatures $T=5$, 10, 50, 100, and 200 K. In order to minimize experimental errors, each of these loops was obtained by averaging over at least five different measurements. The curves cross at $H=0$ where no changes in the resonant frequency are detected, which corresponds to case I in Fig. 4. Additionally, $\Delta\nu$ is also zero at the coercive field (case II). This is a very important feature of this experimental system since it allows measuring the coercivity directly from the experimental curves given in Fig. 5 (a). At first glance we can observe that coercivity does not depend appreciably on temperature, being $\mu_{0}H_{c}\sim 0.1$ T, a result that contrasts strongly with the observed in dense arrays of Ni NWs.
@@
Figure 5: (a) Changes in the resonant frequency, $\Delta\nu$, as a function of the external field, $\mu_{0}H$, for different temperatures as indicated. (b) The corresponding magnetic hysteresis loops calculated using Eqs. (2) and (3). In both figures, the arrows indicate the direction in which $\Delta\nu$ and $\mu_{0}M$ evolve as the external field changes.
@@
As discussed in Sec. II.2, when both the external field and the magnetization are non-zero, and they are parallel to each other, $\Delta\nu>0$, case III in Fig. 4. Instead, if both vectors are antiparallel, $\Delta\nu<0$, see case IV in Fig. 4. This happens when the external field reverses its polarity taking values between 00 and ±$\pm H_{c}$. The positive or negative variations of $\Delta\nu$ can be observed in Fig. 5 (a), noting that the changes in the resonant frequency are, at most, of the order of $1$ Hz.
@@
To calculate the magnetic hysteresis loops we use Eqs. (2) and (3). Our micromechanical torsional oscillator has a resonant frequency close to $56.2$ kHz, a quality factor $Q$ greater than $5\times 10^{4}$, and a moment of inertia along its center rotational axis of $I=(4.0\pm 0.3)\times 10^{-21}$ kg m2. Figure 5 (b) shows the magnetic hysteresis loops that we have obtained from our experimental data. The saturation magnetization does not depend appreciably on temperature and reaches a value of $\mu_{0}M_{s}=(0.6\pm 0.1)$ T, which agrees with the value for bulk Ni, $\mu_{0}M_{b}=0.61$ T Bertotti1998 .
@@
It is interesting to compare one of these magnetic hysteresis loops with an equivalent one (at the same temperature) previously measured for an array of the same Ni NWs still embedded in the alumina template Meneses2018 . Figure 6 shows these curves at $T=5$ K where, for the sake of comparison, we have normalized the magnetization. The differences between them are noticeable. The hysteresis loop of the array displays a S-shape with a low coercive field. This is because the demagnetizing field is not negligible in a dense packing of NWs and, as discussed in the literature Kumar2006 ; Meneses2018 , an additional magnetoelastic effect tends to decrease the effective uniaxial anisotropy constant (and hence to decrease the coercivity). In contrast, the hysteresis loop we have obtained in this work has a more squared shape and the coercive field is appreciably larger. As we have discussed previously, since the interactions between the nanowires are negligible, this curve [and also those shown in Fig. 5 (b)] agrees with the magnetic behavior of a small group of non-interacting Ni NWs.
@@
Figure 6: Hysteresis loops for a dense array of Ni NWs still embedded in the AAO template at $T=5$K (open circles) Meneses2018 , a group of 11 isolated Ni NWs measured with the micromechanical oscillator at the same temperature (solid squares), and a similar single Ni NW measured with a micro-SQUID at $T=6$ K (open squares) Wernsdorfer1996 .
@@
Finally, in Fig. 7 we show the coercive field $\mu_{0}H_{c}$ as function of $T$ for isolated Ni NWs, determined directly from the curves given in Fig. 5 (a). As we observed before qualitatively, this magnitude is nearly constant over the temperature range studied. For comparison, it also shows the coercivity for an array of Ni NWs embedded in the AAO template Meneses2018 , which displays the atypical behavior mentioned throughout this work.
Figure 7: Coercivity versus temperature for a dense array of Ni NWs embedded in the AAO template (open circles) Meneses2018 and the isolated Ni NWs measured with the micromechanical oscillator (solid squares).
@@
where we have considered that $\mu_{0}M_{s}\approx 0.6$ T for an isolated NW and $K_{s}=\mu_{0}M_{s}^{2}/4$, the value of the shape anisotropy constant for an infinitely long rod at saturation. This value is at least twice as large as that measured in our experiments, which corroborates that other relaxation mechanisms are operating in this system.

### v2

@@ Abstract
Magnetic nanowires (NWs) are critical components in fields such as data storage and spintronics, where precise control of their magnetic properties is essential for device optimization. The behavior of isolated NWs is often different from that of an ensemble, offering an opportunity to explore the role that dipolar and magnetoelastic interactions play in the latter system. Unfortunately, the comparison between a collection of NWs and single ones is often poorly characterized, as measuring individual NWs with weak magnetic signals is a challenging task. In this work, we employ a highly sensitive micromechanical torsional oscillator to measure the magnetic response of a few individual Ni NWs with $(72\pm 5)$ nm average diameter, fabricated by electrodeposition in anodic aluminum oxide templates as an array and subsequently released from this membrane. When comparing the magnetic properties as a function of temperature between single NWs and the array, we show that coercivity values of individual NWs are at least twice as large as for the array in the range 55-$200$ K. Also, we characterize the differences in the hysteresis loops, which are more squared for isolated NWs, with a high magnetic remanence $\sim 80\%$ of the saturation value. Our results highlight the crucial role of dipolar and mechanical interactions in modifying the magnetic behavior of NW arrays, providing valuable insights for the design and application of NW-based magnetic devices.
@@ I Introduction
Using the electrodeposition technique, ferromagnetic Ni NWs with a variety of sizes and morphologies can be synthesized inside the pores of an anodic aluminum oxide (AAO) template Razeeb2009 ; Proenca2012 . The magnetic hysteresis properties of these structures have been extensively studied Whitney1993 ; Sellmyer2001 ; Kato2004 ; Kumar2006 ; Meneses2015 ; Meneses2018 ; Campos2020 . In general, within a wide range of temperatures $T$ between $5$ and $300$ K, a monotonic reduction in coercivity $H_{c}$ with increasing $T$ is observed in arrays of Ni NWs with diameters larger than $d\approx 90$ nm which are still embedded in the AAO template Adeela2015 . This is a typical behavior also found in bulk multidomain Ni ferromagnets Hayashi1964 , single-domain Ni nanoparticles Fonseca2003 , and Ni films Hanson1997 . Interestingly, the coercivity of thinner NWs still embedded in the template shows an infrequent behavior: $H_{c}$ increases with increasing $T$ and in some cases it exhibits a maximum around room temperature Kumar2006 ; Meneses2018 ; Campos2020 ; Zeng2002 ; Navas2008 ; Michel2017 .
Authors of previous studies have suggested that the internal microstructure of each wire (i.e., the mean grain size and the magnetocrystalline anisotropy) and the dipolar interactions between the NWs are not the origin of this unexpected behavior in $H_{c}$. Instead, this phenomenon could simply be due to the competition between the shape anisotropy of individual NWs; the magnetoelastic anisotropy arising from the mechanical interaction between the NWs, the AAO template, and the Al substrate; and the dipolar interactions between NWs Kumar2006 ; Meneses2018 ; Zeng2002 . Experimental data seem to confirm these conjectures.
In this work, we use a silicon micromechanical torsional oscillator to study the hysteresis behavior of isolated Ni NWs in a temperature range of 55-$200$ K. In the absence of the template, substrate and dipolar interactions, we measure a coercivity which depends only on the intrinsic characteristics of the NWs.
The outline of the paper is as follows. In Sec. II, we briefly describe how the Ni NWs were synthesized and released from the template, focusing on the most relevant characteristics of the magnetic sample we have studied. We also describe the basic operation of the micromechanical torsional oscillator. In Sec. III, we present the results and discuss their implications. Finally, Sec.IV is devoted to the summary and conclusions.
@@ II Experiments
II.1 Ni NWs
We study Ni NWs that were synthesized in a previous work Meneses2018 . We focus on sample DC65, where the NWs were grown by direct current (DC) electrodeposition inside the pores of an AAO template. In Ref. Meneses2018 , the magnetic hysteresis loops of this sample were measured in a super conducting quantum interference device (SQUID), in a temperature range of 55-$300$ K. The coercivity of sample DC65 increases with increasing $T$ and exhibits a maximum around room temperature. A priori, it is not clear whether this phenomenon reflects the magnetic properties of a single NW or rather the collective behavior of an array of NWs still embedded in the AAO template.
Figure 1: (a) SEM image of the silicon micromechanical oscillator showing the NWs stuck to its plate. This structure is attached to two springs which are in turn anchored to the substrate, allowing torsional oscillations forced by an alternating voltage signal connected to the electrodes. When a magnetic field $\mathbf{H}$ is applied, the resonance frequency of the device changes, allowing us to infer the magnetic moment of the sample. (b) SEM image of a typical Ni NW.
@@ II.1 Ni NWs
[…]
More detailed SEM images show that some of the fragments (supposedly individual NWs) that were glued on the microsensor are bundles formed from just a few NWs, as displayed in Fig. 2. Since these clusters are made up of a few NWs, the dipolar interactions between them are extremely low. To estimate the order of magnitude of this unwanted contribution, we roughly calculate the longitudinal component of the magnetic field $H_{z}$ produced by a NW at a transverse distance $r$ from its center of mass, describing this object by two magnetic charges separated by a distance $l$ (the length of the NW) Kumar2006 ; Bertotti1998 :
@@
[…]
@@ II.2 Micromagnetometer
The silicon micromechanical torsional oscillator was manufactured in the MEMSCAP Inc. Foundry MEMSCAP . Similar microdevices have been previously used as micromagnetometers of high sensitivity to measure the hysteresis behavior of manganite nanotubes Dolz2008 ; Antonio2010 ; Dolz2020 as well as mesoscopic samples of a high-$T_{c}$ superconductor Dolz2007 ; Dolz2010 . In simple terms, its operation can be understood as follows. The whole system (the micromechanical oscillator with the NWs stuck on its plate) is cooled under vacuum inside a helium closed-cycle cryogenerator until reaching a given temperature. Then the microdevice is actuated electrostatically by means of a function generator, and its movement is sensed capacitively using a lock-in amplifier. In this way, the resonance frequency of the system is accurately measured which depends, among other things, on the elastic constant of its serpentine springs. When a uniform magnetic field (provided by an electromagnet) is applied along the easy axis of the NWs [see Fig. 1 (a)], the plate of the oscillator experiences an additional torque, and therefore, the resonance frequency changes. Although extremely small, these shifts in resonance frequency (<1<1 Hz) can be measured and used to infer the magnetization state of the sample (see below). Additional details of the experimental setup can be found in Refs. Dolz2008 ; Antonio2010 .
@@
Figure 3: Resonance curves obtained when $\mathbf{H}$ or $\mathbf{M}$ (or both) are zero, and therefore, there is no additional magnetic torque (cases I and II), and when these fields are parallel (case III) or antiparallel (case IV), which results in the appearance of a restoring or antirestoring magnetic torque, respectively.
@@
Typical resonance curves are plotted in Fig. 3. With red squares, we show the amplitude of oscillation as a function of frequency when there is no additional magnetic torque on the system. This happens when either no external field $\mathbf{H}$ is applied or when the magnetization $\mathbf{M}$ of the sample is zero, which correspond, respectively, to cases I and II drawn in Fig. 4 (furthermore, there is no magnetic torque in the trivial case where both vectors are zero). When performing a hysteresis experiment, case II arises at the coercive field $\mathbf{H}_{c}$ since, in this instance, $\mathbf{M}=0$.
@@
Figure 4: Sketches corresponding to cases I, II, III, and IV. No additional magnetic torque arises when (a) $\mathbf{H}=0$ or (b) $\mathbf{M}=0$ . However, when both fields are nonzero, (c) a restoring torque appears if $\mathbf{H}$ and $\mathbf{M}$ are parallel, while (d) this torque is antirestoring if $\mathbf{H}$ and $\mathbf{M}$ are antiparallel.
@@
A different situation appears when both $\mathbf{H}$ and $\mathbf{M}$ are nonzero. As shown in case III in Fig. 4, if these vectors are parallel, any slight twist with respect to the equilibrium position results in the appearance of a restoring magnetic torque, which adds to the torque produced by the serpentine springs. As a consequence, the resonance frequency must increase (note that, in any mechanical oscillator, as the restoring force or torque increases, the resonant frequency also increases). The corresponding curve in Fig. 3 (blue circles) shows this positive shift in frequency. On the contrary, if $\mathbf{H}$ and $\mathbf{M}$ are antiparallel as in case IV, the sample experiences an antirestoring magnetic torque, which reduces the total torque applied to the oscillator plate, and then the resonance frequency decreases, see the corresponding resonant curve in Fig. 3 (magenta triangles).
@@
The above qualitative description can be transformed into a quantitative formalism by making a series of simple analytical calculations and well-founded approximations. Technical details are given in Appendix A. Such calculations lead to a simple expression that allows the magnetization of the sample to be estimated from the changes in the resonance frequency $\Delta\nu$ measured in our experiments:
@@
where $\nu_{0}$ is the natural resonance frequency of the micromechanical torsional oscillator (when the magnetic torque is zero), $I$ is the moment of inertia of the whole system along its center rotational axis, and as we defined above, $V_{n}$ is the volume of the magnetic sample.
@@ III Results and discussion
Figure 5(a) shows the hysteresis loops of $\Delta\nu$ vs $H$ measured at temperatures $T=5$, 10, 50, 100, and 200 K. To minimize experimental errors, each of these loops was obtained by averaging over at least five different measurements. The curves cross at $H=0$ where no changes in the resonant frequency are detected, which corresponds to case I in Fig. 4. Additionally, $\Delta\nu$ is zero at the coercive field (case II). This is a very important feature of this experimental system since it allows measuring the coercivity directly from the experimental curves given in Fig. 5(a). At first glance we can observe that coercivity does not depend appreciably on temperature, being $\mu_{0}H_{c}\sim 0.1$ T, a result that contrasts strongly with the observed in dense arrays of Ni NWs.
@@
Figure 5: (a) Changes in the resonant frequency $\Delta\nu$ as a function of the external field $\mu_{0}H$ for different temperatures, as indicated. (b) Corresponding magnetic hysteresis loops calculated using Eqs. (2) and (3). In both figures, the arrows indicate the direction in which $\Delta\nu$ and $\mu_{0}M$ evolve as the external field changes.
@@
As discussed in Sec. II.2, when both the external field and the magnetization are nonzero and they are parallel to each other, $\Delta\nu>0$, case III in Fig. 4. Instead, if both vectors are antiparallel, $\Delta\nu<0$, see case IV in Fig. 4. This happens when the external field reverses its polarity, taking values between 0 and $\pm H_{c}$. The positive or negative variations of $\Delta\nu$ can be observed in Fig. 5(a), noting that the changes in the resonant frequency are, at most, of the order of $1$ Hz.
@@
To calculate the magnetic hysteresis loops we use Eqs. (2) and (3). Our micromechanical torsional oscillator has a resonant frequency $\nu_{0}\sim 56.2$ kHz, a quality factor $Q$ that roughly ranges from $1.7\times 10^{5}$ to $1.5\times 10^{4}$, and a moment of inertia along its center rotational axis of $I=(4.0\pm 0.3)\times 10^{-21}$ kg m2. In Appendix B, we show how $\nu_{0}$, $Q$, and the sensitivity of this microdevice depend on temperature.
Figure 5(b) shows the magnetic hysteresis loops that we have obtained from our experimental data. The saturation magnetization does not depend appreciably on temperature and reaches a value of $\mu_{0}M_{s}=(0.6\pm 0.1)$ T, which agrees with the value for bulk Ni, $\mu_{0}M_{b}=0.61$ T Bertotti1998 .
@@
It is interesting to compare one of these magnetic hysteresis loops with an equivalent one (at the same temperature) previously measured for an array of the same Ni NWs still embedded in the alumina template Meneses2018 . Figure 6 shows these curves at $T=5$ K where, for the sake of comparison, we have normalized the magnetization. The differences between them are noticeable. The hysteresis loop of the array displays an S-shape with a low coercive field. This is because the demagnetizing field is not negligible in a dense packing of NWs and, as discussed in the literature Kumar2006 ; Meneses2018 , an additional magnetoelastic effect tends to decrease the effective uniaxial anisotropy constant (and hence to decrease the coercivity). In contrast, the hysteresis loop we have obtained in this work has a more squared shape, and the coercive field is appreciably larger. As we have discussed previously, since the interactions between the NWs are negligible, this curve [and those shown in Fig. 5(b)] agrees with the magnetic behavior of a small group of noninteracting Ni NWs.
@@
Figure 6: Hysteresis loops for a dense array of Ni NWs still embedded in the AAO template at T=5T=5K (open circles) Meneses2018 , a group of $1$ isolated Ni NWs measured with the micromechanical oscillator at the same temperature (solid squares), and a similar single Ni NW measured with a micro-SQUID at $T=6$ K (open squares) Wernsdorfer1996 .
@@
Finally, in Fig. 7 we show the coercive field $\mu_{0}H_{c}$ as a function of $T$ for isolated Ni NWs, determined directly from the curves given in Fig. 5(a). As we observed before qualitatively, this magnitude is nearly constant over the temperature range studied. For comparison, it also shows the coercivity for an array of Ni NWs embedded in the AAO template Meneses2018 , which displays the atypical behavior mentioned throughout this work.
Figure 7: Coercivity vs temperature for a dense array of Ni NWs embedded in the AAO template (open circles) Meneses2018 and the isolated Ni NWs measured with the micromechanical oscillator (solid squares).
@@
where we have considered that $\mu_{0}M_{s}\approx 0.6$ T for an isolated NW and $K_{s}=\mu_{0}M_{s}^{2}/4$, the value of the shape anisotropy constant for an infinitely long rod at saturation. This value is at least twice as large as that measured in our experiments, which corroborates that other relaxation mechanisms are operating in this system. In this regard, numerical calculations show that for small diameter Ni NWs, such as those studied in this work, the magnetization is reversed by the propagation of a Bloch-type domain wall Hertel2001 .
One approach to experimentally analyze the switching mechanisms is to study how the coercivity varies with the angle at which the external magnetic field is applied Wernsdorfer1996 . Unfortunately, our experimental setup makes it very difficult to explore this problem in greater depth. Although, in theory, such a study could be conducted using our microdevice Kamra2014 , applying the external field within the plane of the oscillator plate causes the torque component perpendicular to the oscillation plane to decrease as the angle of application increases. Consequently, the resulting change in the resonant frequency is significantly reduced (in fact, this torque component becomes zero when the angle reaches $90^{\circ}$). In practice, this means that coercivity can only be measured over a very limited angular range, making it challenging to accurately determine the magnetization reversal mechanism.